OpenStudy (anonymous):

A triangle has an area of 4, a base of 2x+4, and a height of x-1. Solve for x and provide a justification for each step.

OpenStudy (anonymous):

(2x + 4)(x -1) / 2 = 4 (2x + 4)(x -1) = 8 x = 2 or -3

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

4 = (2x + 4)(x - 1)/2 8 = 2x^2 - 2x + 4x - 4 2x^2 + 2x - 12 = 0 2(x^2 + x - 6) = 0 (x + 3)(x - 2) = 0 x + 3 = 0 or x - 2 = 0 x = -3 or x = 2